If $10,000 is invested at an interest rate of 2% per year, compounded semiannually, find the value of the investment after the given number of years. (Round your answers to the nearest cent.)
a)6 years
b)12 years
c)18 years

Answer by lwsshak3(11628)

If $10,000 is invested at an interest rate of 2% per year, compounded semiannually, find the value of the investment after the given number of years. (Round your answers to the nearest cent.)
a)6 years
b)12 years
c)18 years
***
compound interest formula: A=P(1+r/n)^nt, P=initial investment, r=interest rate, n=number of compounding periods per year, A=amt after t-years.
For given problem:
P=10000
r=.02
n=2
t=6, 12, 18
..
A(6)=10000(1+.02/2)^2*6
A(6)=10000(1.01)^12=11,268.25
A(12)=10000(1.01)^24=12,697.35
A(18)=10000(1.01)^36=14,307.69